OpenStudy (anonymous):

summation of n=1 to infinty ln((n^2+1)/(2n^2+2)) Does this diverge or converge?

OpenStudy (anonymous):

it converges.

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \ln ((n ^{2}+1)/(2n ^{2}+2))\]

OpenStudy (anonymous):

how does this converge?

OpenStudy (anonymous):

ln(n^2+1)/(2(n^2+1)) = ln 1/2 is a finite number

OpenStudy (anonymous):

Yes, but for every n you add ln(1/2) And there are a lot of ns...

OpenStudy (anonymous):

sorry i had this one already. my question is rather this: \[\sum_{1}^{\infty} \ln ((n ^{2}+1)/(n ^{2}+2))\]

OpenStudy (anonymous):

wait, I see you posted the problem again in equation form. Let me see

OpenStudy (anonymous):

yes. the denominator in this case is n squared plus 2

OpenStudy (anonymous):

Newton, I have a doubt, \[\sum_{n=1}^{\infty}1/2 = ?\]

OpenStudy (anonymous):

with the first one, the limit of the expression was 1/2 and the natural log was not equal to zero so it indeed converged. but i can't determine this one.

OpenStudy (anonymous):

that wouldn't make any sense since there is no n term in the summation. Isn't that right?

OpenStudy (anonymous):

It isn't summing a half once though, it is summing it for every n. n = 1 , add ln(1/2) n = 2 , add ln(1/2) n = .... etc

OpenStudy (anonymous):

okay i understand.

OpenStudy (anonymous):

do you guys have a clue for the question?

OpenStudy (anonymous):

or should repost the question again? if you look above, you'd realize i posted a different but quite similar question.

OpenStudy (anonymous):

for the first problem, it is infinite, for the second problem, as n reaches infinity, the summation reaches 0. So it is finite

OpenStudy (anonymous):

how? that's what i am not getting.

OpenStudy (anonymous):

not the summation, the value of n^2+1/n^2/2

OpenStudy (anonymous):

it's n^2+1/n^2+2

OpenStudy (anonymous):

oh right, my mistake. Let me start afresh.

OpenStudy (anonymous):

i tried the divergence test and it doesn't work. does it converge or diverge?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(n ^{2}+1)/(n ^{2}+2)\]

OpenStudy (anonymous):

that is the problem correct?

OpenStudy (anonymous):

whoops, its the natural log of that

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

we know that ln 1 =0

OpenStudy (anonymous):

for large values of n, \[n ^{2}+1/n ^{2}+2 \approx n ^{2}/n ^{2} =1\]

OpenStudy (anonymous):

so as n approaches infinity, the natural log of the given equation approaches 0. so it is a finite number

OpenStudy (anonymous):

therefore it converges

OpenStudy (anonymous):

what i did with the first question was to find the limit which was 1/2 and apply the natural log. With this question, what was your approach?

OpenStudy (anonymous):

In general, the geometric series \[\sum_{n=0}^{\infty}z ^{n}\] converges if and only if |z| < 1.

OpenStudy (anonymous):

so you used the absolute convergence test i presume, which i am yet to learn. is that the only way out? can you outline the specifics of your step?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \ln((n^2+1)/(2n^2+2))\] this sum does not converge

OpenStudy (anonymous):

i think you're right since the ln 1/2 is not equal to zero for the divergence test, hence diverges.

OpenStudy (anonymous):

do you see the other question?

OpenStudy (anonymous):

what for example?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \ln ((n ^{2}+1)/(n ^{2}+2))\]

OpenStudy (anonymous):

this sum we can find

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

there are two ways to approach this problem. I am not sure which way is correct though. So here goes. \[(n ^{2}+1)/(n ^{2}+2) = (n ^{2}+2-1)/(n ^{2}+2)\]

OpenStudy (anonymous):

dhatraditya you right

OpenStudy (anonymous):

\[=1-1/(n ^{2}+2)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but we have natural logarithm

OpenStudy (anonymous):

for large values of n, 1/n^2+2 approaches 0. so for large values of n, 1- 1/n^2+2 = 1-0 = 1 ln 1 = 0

OpenStudy (anonymous):

no it's not good i think we must do like as...

OpenStudy (anonymous):

yes, ln 1 = 0 even for natural log. log 1 to any base is 0.

OpenStudy (anonymous):

\[\sum_{1}^{\infty}\ln(n^2+1)-\sum_{1}^{\infty}\ln(n^2+2)\]

OpenStudy (anonymous):

you guys have me confused now. which approach is agreed upon now. and does it converge or not? can you also state the test you are using?

OpenStudy (anonymous):

it converges. That is agreed upon. The method is what we are disagreeing upon.

OpenStudy (anonymous):

what i have learnt so far uses the various test methods like divergence test, integral test, basic comparison test, etc to solve this questions. which one are you applying in this case?

OpenStudy (anonymous):

but this sums does not converge =)

OpenStudy (anonymous):

are you talking about his sum or the previous one, hobbit?

OpenStudy (anonymous):

because as far as I can tell, this sum seems to be converging.

OpenStudy (anonymous):

but this sums does not converge =)

OpenStudy (anonymous):

sorry i have some problems with charge on my mac)))

OpenStudy (anonymous):

oh well, then you should explain it to bbcram. because I seem to be wrong on this one.

OpenStudy (anonymous):

what do you don't understand bbcram?

OpenStudy (anonymous):

so guys i have a question)) where you from?

OpenStudy (anonymous):

im from india Please explain how you got this sum is diverging.

OpenStudy (anonymous):

cool i am from Russia you must to know that ln(a/b)=ln(a)-ln(b)

OpenStudy (anonymous):

Russian eh? Cool. Student? yes ln(a/b)=lna-lnb

OpenStudy (anonymous):

and \[\sum_{?}^{?}(a-b)=\sum_{?}^{?}a-\sum_{?}^{?}b\]

OpenStudy (anonymous):

Does Hobbit think: \[\sum_{1}^{\infty} \ln \frac{n^2+1}{n^2+2}\] DIVERGES? xD

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes, he seems to think so. I am asking him why he thinks that

OpenStudy (anonymous):

answer near -0.956448

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

don't worry Hobbit, Wolfram |Alpha thinks so too ... xD

OpenStudy (anonymous):

That's not divergence, kiddo.

OpenStudy (anonymous):

it s partial sum

OpenStudy (anonymous):

INewton where you from? and where do you study?

OpenStudy (anonymous):

sorry guys my pc suddenly restarted. It's no less clear now. INewton, dhatraditya, Hobbit. I am in the united states. I just started the topic on sequence and series this week and were challenged with this question. I have worked on others successfully but this is a headache. So what is your take? I seem to be seeing conflicting answers and approaches

OpenStudy (anonymous):

that is a finite number. So it is convergent.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then why did you think it was divergent?

OpenStudy (anonymous):

i dont think that it is divergent

OpenStudy (anonymous):

you said it does not converge.

OpenStudy (anonymous):

∑1∞ln(n2+1) is divergent and ∑1∞ln(n2+2) is divergent

OpenStudy (anonymous):

is not it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok, so the ln of either the numerator or denominator in this case diverges. cool. however, the ln for the equation is 0 right? which makes it convergent?

OpenStudy (anonymous):

i don't know, but may be gamma function ?

OpenStudy (anonymous):

i in confuse too now =)

OpenStudy (anonymous):

i think dhatraditya is right.

OpenStudy (anonymous):

@dhatradtiya, can you just once again write the steps you took so i can write that down and study your procedure?

OpenStudy (anonymous):

yes, \[(n ^{2}+1)/(n ^{2}+2)=((n ^{2}+2)−1))/(n ^{2}+2)\]

OpenStudy (anonymous):

\[=1−1/(n ^{2}+2)\]

OpenStudy (anonymous):

for large values of n,

OpenStudy (anonymous):

\[1/(n ^{2}+2) \approx0\]

OpenStudy (anonymous):

yes and so

OpenStudy (anonymous):

therefore \[1−1/(n ^{2}+2) \approx1-0 = 1\]

OpenStudy (anonymous):

but 1-1=0 not 1

OpenStudy (anonymous):

\[n \rightarrow \infty, \ln (1-1/(n ^{2}+2) = 0\]

OpenStudy (anonymous):

for n=1 ln(1-1/3) for n=2 ln(1-1/6) for n=3 ln(1-1/11) for n=infty ln(1-0)=0

OpenStudy (anonymous):

yes but we must to find sum of first members

OpenStudy (anonymous):

ln(1-1/3)+ln(1-1/6)+ln(1-1/11)+...